# RelaCtrl: Relevance-Guided Efficient Control for Diffusion Transformers

### v4

@@ 3.1 Experiment Setup
Baseline. We conducted a comparative analysis of RelaCtrl against state-of-the-art (SOTA) control techniques, including Uni-ControlNet (Zhao et al., 2024), UniControl (Qin et al., 2023), ControlNet-XS (Zavadski et al., 2025), ControlNext (Peng et al., 2024), and PixArt-$\delta$ (Chen et al., 2024). For the first four methods, we utilized their officially released pre-trained weights. It is worth noting that ControlNet-XS and ControlNext only provided weights for Canny and Depth conditions, with the Canny weights for ControlNext being trained on an animation dataset. To further highlight the performance of our proposed method, we trained the control branch for both PixArt-$\delta$ and our method entirely from scratch under identical experimental settings, enabling a rigorous quantitative and qualitative comparison.
3.2 Compared with SOTA methods
@@
Fig. 4 presents a visual comparison of RelaCtrl and other methods under Canny and Depth conditions. Our method significantly reduces computational complexity and parameter usage while maintaining generation quality comparable to PixArt-ControlNet, outperforming other approaches based on SD1.5 and SDXL. Additional controllable generation results of RelaCtrl are shown in Fig. 5, further illustrating its effectiveness in extracting and injecting information from conditional images during the generation process. This enables the production of results that align seamlessly with conditional controls.
Figure 4: Qualitative comparison of different methods. Please zoom in for better details.
@@
3.3 Algorithm efficiency analysis
@@
We conducted a comprehensive analysis of RelaCtrl and ControlNet with 13 replicated modules (PixArt-$\delta$) under consistent experimental settings. Specifically, the input data was in bfloat16 format with a resolution of 512, the test batch size was set to 1, and all experiments were performed on an NVIDIA A100 GPU. Table 2 presents a comparative evaluation of the two methods, with each indicator expressed as a percentage relative to the original base model. Notably, memory consumption excludes usage by the CLIP and T5 encoders. Compared to the original PixArt-$\alpha$, RelaCtrl incurred a modest 7.38% increase in parameters and an 8.61% increase in computational complexity. These increments are significantly lower than those of the ControlNet method, which demonstrates nearly a 50% increase in both parameters and complexity. Additionally, RelaCtrl showed some advantages in memory usage and inference time, although it is worth noting that inference time is predominantly influenced by the speed of the main network rather than the control branch, leading to similar small increases in this metric for both methods. Overall, RelaCtrl outperforms the original ControlNet method applied to the DiT model by achieving comparable or superior metrics and visual results while significantly reducing computational resource consumption. This demonstrates the effectiveness and efficiency of the proposed RelaCtrl framework.
@@
Table 2: Evaluation of the proposed method’s effectiveness, with the following units for the four metrics: Parameters (M), Complexity (GFLOPs), Inference Time (s), and Memory Usage (MiB)
Method Parameters Complexity Inference Memory
PixArt-$\alpha$ 611.15 542.56 3.81 2114
w/ ControlNet +294.34 +270.57 +0.51 +1694
(+48.16%) (+49.87%) (+13.39%) (+80.13%)
w/ RelaCtrl +45.15 +46.71 +0.24 +395
(+7.38%) (+8.61%) (+6.30%) (+18.70%)
3.4 Ablation study
@@
We conducted quantitative experiments to evaluate the efficacy of guiding control block placement based on DiT-ControlNet Relevance. Utilizing relevance scores, we ranked the control block positions from highest to lowest and placed copy blocks at the top-ranked positions, evaluating configurations with 13, 12, 11, and 10 blocks. The results are summarized in Table 3. As anticipated, reducing the number of control blocks resulted in a gradual decline in FID and HDD metrics. Notably, under the guidance of relevance scores, using only 11 control blocks achieved performance comparable to the original ControlNet with 13 replicated blocks. When further reduced to 10 blocks, the quality of the generated results slightly declined relative to the 13-block setup. Consequently, we selected the top 11 positions ranked by relevance scores for control placement, as illustrated by the white numbers in Fig. 2.
@@
To further validate the significance of the RGLC block and the incorporation of Prior 2 based on revelation, we conducted additional quantitative experiments which are shown in Table 4. Employing the ControlNet with 13 copy blocks as the baseline and the RelaCtrl network as the foundation, we evaluated the impact of removing the RGLC block and relevance prior 2, respectively. The results demonstrate that the absence of either component leads to a decline in image quality and control performance. Replacing the RGLC block with the original copy block not only significantly increases the parameter scale but also results in performance degradation, highlighting the efficacy of the RGLC block. In the experiment (w/o Prior 2), we applied a uniform TDSM channel division across all RGLC blocks and eliminated our correlation-specific settings. The poorer HDD and FID metrics observed indicate that allocating more parameters and computational resources to control positions with higher correlation is crucial for achieving high-quality controllable generation. These results underscore the importance of both the RGLC block and relevance prior 2 in enhancing generation quality and control precision.
@@
Table 4: The impact of the RGLC block and the number of TDSM partitions within it on generation performance. The PixArt-$\delta$ with 13 coped blocks serves as the baseline for parameter comparison.
@@
3.5 RelaCtrl in Flux
To further validate the effectiveness and generalizability of the proposed method, we conducted additional experiments on Flux.1-dev, with the results provided in the Appendix D.2
4 Conclusion
@@
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Related works
A.1 Diffusion-based models
In recent years, diffusion-based methods have garnered significant success in the field of generation (He et al., 2024; Feng et al., 2024; Zhang et al., 2024), particularly in text-to-image (T2I) generation (Guo et al., 2023; Shuai et al., 2024). These methods utilize text embeddings derived from pre-trained language encoders such as CLIP (Radford et al., 2021), Bert (Devlin, 2018) and T5 (Raffel et al., 2020) to generate images with high fidelity and diversity through an iterative denoising process. Recently, the introduction of the latent diffusion model(Rombach et al., 2022) has marked a significant advancement in this field, enhancing the quality and efficiency of generated content. In pursuit of greater scalability and enhanced generation quality, models like DiT (Peebles & Xie, 2023; Chen et al., 2023)integrate large-scale Transformer architectures into the diffusion framework, pushing the boundaries of generative performance. Building on this foundation, Flux (BlackForestlabs AI, 2024) synthesizes flow-matching (Lipman et al., 2022) and Transformer-based architecture to achieve state-of-the-art performance.
A.2 Controllable generation with diffusion models
Controllable generation has emerged as a prominent area of research in diffusion models (Zhang et al., 2023; Qin et al., 2023; Zhao et al., 2024; Chen et al., 2024; Peng et al., 2024). Currently, there are two main approaches to incorporating controllable conditions into image generation: (1) training a large diffusion model from scratch to enable control under multiple conditions, and (2) fine-tuning a lightweight structure while keeping the original pre-trained model frozen. However, the first approach demands significant computational resources, which limits its accessibility for broader dissemination and personal use. In contrast, recent studies have explored the addition of supplementary network structures to pre-trained diffusion models, allowing for control over the generated outputs without the need to retrain the entire model.
ControlNet (Zhang et al., 2023) enables image generation that aligns with control information by reproducing specific layers within the network and connecting them to the original layers using zero convolution. Building on this foundation, Uni-Control (Qin et al., 2023) introduces a Mixture-of-Experts (MoE) framework, which unifies control across multiple spatial conditions. ControlNet-XS (Zavadski et al., 2025) further improves the interaction bandwidth and frequency between the control and main branches within the ControlNet architecture, drawing inspiration from principles of feedback control systems. Nonetheless, these approaches are primarily based on the U-Net structure and may not yield the desired results when directly applied to Diffusion Transformers (DiT) without modification (Chen et al., 2024). PixArt-$\delta$ (Chen et al., 2024) proposed a design methodology specifically tailored for DiT, but directly copying the first half of the network results in a 50% increase in both parameter count and computational complexity, resulting in high computational cost and inconvenient for community research and practical deployment.
Appendix B Proof to Theorem
B.1 Proof to Theorem 3.4
@@ Appendix B Proof to Theorem
$\begin{split}\sum\limits_{w=0}^{W{{d}_{i}}-1}{\left|w-{{t}_{wj}}\right|}&=\sum\limits_{w=0}^{{{t}_{hj}}}{\left|{{t}_{wj}}-w\right|}+\sum\limits_{w={{t}_{hj}}}^{W{{d}_{i}}-1}{\left|w-{{t}_{wj}}\right|}\\ &=\frac{{{t}_{wj}}({{t}_{wj}}+1)}{2}+\frac{(W{{d}_{i}}-{{t}_{wj}})(W{{d}_{i}}-{{t}_{wj}}-1)}{2}\end{split}$ (10)
@@
$\begin{split}d({{t}_{j}})&\geq\frac{\sqrt{2}}{4(HW{{d}_{i}}-1)}[H{{m}_{wj}}({{m}_{wj}}+1)+W{{d}_{i}}{{m}_{hj}}({{m}_{hj}}+1)\\ &+H(W{{d}_{i}}-{{m}_{wj}})(W{{d}_{i}}-{{m}_{wj}}-1)+W{{d}_{i}}(H-{{m}_{hj}})(H-{{m}_{hj}}-1)]\end{split}$ (11)
@@
Among these, $d({{t}_{j}}\left|HW{{d}_{i}}-j\right.)$ represents the expected interactive distance when the number of remaining tokens is $HW{{d}_{i}}-j$. Based on Assumption 2.1, we know $j\in[2,{{s}^{2}}-1]$, and we can derive the following approximation:
@@
Appendix C Training Details
Training and Testing Datasets. We curated a dataset containing 1.73 million high-resolution, high-quality images, each with an aesthetic score of 5.5 or higher. Following the production methodology outlined in ControlNet (Zhang et al., 2023), we generated various types of conditional images, including HED, Canny, Depth, and Segment. For the quantitative experiments, the model was trained at a resolution of 512 and evaluated on the COCO validation dataset (Lin et al., 2014), consisting of 5000 images. For qualitative visual evaluation, the model was trained at a resolution of 1024 and tested on a high-quality, high-resolution test set of 1000 images. To ensure a fair evaluation, all experiments were conducted with a training of 5 epochs on 16 NVIDIA A100 GPUs, using identical settings across all trials.
Appendix D Additional experimental results
D.1 More exploration on Relevance
@@
D.2 Experiment on Flux
@@
In this section, we apply RelaCtrl to the latest Flux.1-dev with 12 billion parameters for experimentation. The comparison methods include OminiControl (Tan et al., 2024) and Flux ControlNet which is implemented based on the Diffusers library. Flux ControlNet uses the standard configuration of a double-stream layer with the first four replication layers, alongside a single-stream layer with the first ten replication layers.
When applying RelaCtrl to Flux, we aim to leverage as many insights and designs from the main text as possible. Accordingly, we retain the control positions of the double-stream layer unchanged and select seven positions out of ten on the single-stream layer for control. Table 5 presents the efficiency comparison between the methods, with GFLOPs calculated at a resolution of 512, and inference time measured at 512 resolution with 30 DDIM sample steps. Although OminiControl introduces only a modest increase in parameters, it leads to a significant rise in computational load and reasoning time, severely limiting its practical efficiency. Flux ControlNet, on the other hand, shows moderate increases across several efficiency metrics. In contrast, RelaCtrl for Flux significantly reduces both computational complexity and reasoning time, achieving an optimal balance between efficiency and parameter volume. Fig. 8 illustrates the visual results of various control methods on Flux.1-dev. Among them, RelaCtrl not only achieves precise and detailed control effects but also offers significant advantages in computational efficiency, demonstrating the applicability and generalization of the proposed method across different DiT architectures.
Table 5: Effectiveness comparison experiment conducted on Flux.1-dev model.
Setting Parameters (M) Complexity (GFLOPs) Inference (s)
Flux.1-dev 11901.39 9925.78 4.78
+OminiControl +14.49 +6637.41 +3.11
+ControlNet +2952.65 +2578.33 +0.79
+RelaCtrl +549.19 +495.03 +0.34
Figure 8: Visual comparison of different control methods on Flux.1-dev.
D.3 Inference with Community Models
@@
We perform inference using the PixArt weights which were fine-tuned with Lora. Although RelaCtrl has not previously trained in these weights, it can still utilize them effectively. Fig. 9 showcases the model’s generated paint, oil, gufeng, and pixel-style images under the specified conditions.

### v5

@@ Methods
Baseline. We conducted a comparative analysis of RelaCtrl against state-of-the-art (SOTA) control techniques, including Uni-ControlNet (Zhao et al. 2024), UniControl (Qin et al. 2023), ControlNet-XS (Zavadski et al. 2025), ControlNext (Peng et al. 2024), and PixArt-$\delta$ (Chen et al. 2024). For the first four methods, we utilized their officially released pre-trained weights. It is worth noting that ControlNet-XS and ControlNext only provided weights for Canny and Depth conditions, with the Canny weights for ControlNext being trained on an animation dataset. To further highlight the performance of our proposed method, we trained the control branch for both PixArt-$\delta$ and our method entirely from scratch under identical experimental settings, enabling a rigorous quantitative and qualitative comparison.
Figure 3: Qualitative comparison of different methods. From left to right, each row shows the conditioning input, UniControl, Uni-ControlNet, ControlNet-XS, ControlNext, PixArt-$\delta$, and our RelaCtrl.
Compared with SOTA Methods
@@
Fig. 3 presents visual comparisons of RelaCtrl and other methods under Canny and Depth conditions. Our method significantly reduces computational complexity and parameter usage while maintaining generation quality comparable to PixArt-ControlNet, outperforming other approaches. Additional controllable generation results of RelaCtrl are shown in Fig. 4, further illustrating its effectiveness in extracting and injecting information from conditional images during the generation process. This enables the production of results that align seamlessly with conditional controls.
@@
Algorithm Efficiency Analysis
@@
We conducted a comprehensive analysis of RelaCtrl and ControlNet with 13 replicated modules (PixArt-$\delta$) under consistent experimental settings. Specifically, the input data was in bfloat16 format with a resolution of 512, the test batch size was set to 1, and all experiments were performed on an NVIDIA A100 GPU. Table 2 presents a comparative evaluation of the two methods, with each indicator expressed as a percentage relative to the original base model. Notably, memory consumption excludes usage by the CLIP and T5 encoders. Compared to the original PixArt-$\alpha$, RelaCtrl incurred a modest 7.38% increase in parameters and an 8.61% increase in computational complexity. These increments are significantly lower than those of the ControlNet method, which demonstrates nearly a 50% increase in both parameters and complexity. Additionally, RelaCtrl showed some advantages in inference time, although it is worth noting that inference time is predominantly influenced by the speed of the main network rather than the control branch, leading to similar small increases in this metric for both methods. Overall, RelaCtrl outperforms the original ControlNet method applied to the DiT model by achieving comparable or superior metrics and visual results while significantly reducing computational resource consumption. This demonstrates the effectiveness and efficiency of the proposed RelaCtrl framework.
Method Parameters Complexity Inference
PixArt-$\alpha$ 611.15 542.56 3.81
w/ ControlNet +294.34 +270.57 +0.51
(+48.16%) (+49.87%) (+13.39%)
w/ RelaCtrl +45.15 +46.71 +0.24
(+7.38%) (+8.61%) (+6.30%)
@@
Table 2: Evaluation of the proposed method’s effectiveness, with the following units for the three metrics: Parameters (M), Complexity (GFLOPs), Inference Time (s).
Ablation Study
@@
We conducted quantitative experiments to evaluate the efficacy of control block placement based on ControlNet Relevance. Utilizing relevance scores, we ranked the control block positions from highest to lowest and placed copy blocks at the top-ranked positions, evaluating configurations with 13, 12, 11, and 10 blocks. The results are summarized in Table 3. As anticipated, reducing the number of control blocks resulted in a gradual decline in FID and HDD metrics. Notably, under the guidance of relevance scores, using only 11 control blocks achieved performance comparable to the original ControlNet with 13 replicated blocks. When further reduced to 10 blocks, the quality of the generated results slightly declined relative to the 13-block setup. Consequently, we selected the top 11 positions ranked by relevance scores for control placement.
@@
To further validate the significance of the RGLC block and the incorporation of Prior 2 based on revelation, we conducted additional quantitative experiments, which are shown in Table 4. Employing the ControlNet with 13 copy blocks as the baseline and the RelaCtrl network as the foundation, we evaluated the impact of removing the RGLC block and relevance prior 2, respectively. The results demonstrate that the absence of either component leads to a decline in image quality and control performance. Replacing the RGLC block with the original copy block not only significantly increases the parameter scale but also results in performance degradation, highlighting the efficacy of the RGLC block. In the experiment (w/o Prior 2), we applied a uniform TDSM channel division across all RGLC blocks and eliminated our correlation-specific settings. The poorer HDD and FID metrics observed indicate that allocating more parameters and computational resources to control positions with higher correlation is crucial for achieving high-quality controllable generation. These results underscore the importance of both the RGLC block and relevance prior 2 in enhancing generation quality and control precision.
@@
Table 4: The impact of the RGLC block and the number of TDSM partitions within it on generation performance. The PixArt-$\delta$ with 13 copied blocks serves as the baseline for parameter comparison.
Conclusion
@@
Acknowledgments
This work was supported by the Hefei Municipal Natural Science Foundation, HZR2417.
Supplementary Material
Appendix A Proof to Theorem
Proof of Theorem 4
@@ Appendix A Proof to Theorem
$\begin{split}\sum\limits_{w=0}^{W{{d}_{i}}-1}{\left|w-{{t}_{wj}}\right|}&=\sum\limits_{w=0}^{{{t}_{hj}}}{\left|{{t}_{wj}}-w\right|}+\sum\limits_{w={{t}_{hj}}}^{W{{d}_{i}}-1}{\left|w-{{t}_{wj}}\right|}\\ &=\frac{{{t}_{wj}}({{t}_{wj}}+1)}{2}\\ &+\frac{(W{{d}_{i}}-{{t}_{wj}})(W{{d}_{i}}-{{t}_{wj}}-1)}{2}\end{split}$ (10)
@@
$\begin{split}d({{t}_{j}})&\geq\frac{\sqrt{2}}{4(HW{{d}_{i}}-1)}[H{{m}_{wj}}({{m}_{wj}}+1)\\ &+W{{d}_{i}}{{m}_{hj}}({{m}_{hj}}+1)\\ &+H(W{{d}_{i}}-{{m}_{wj}})(W{{d}_{i}}-{{m}_{wj}}-1)\\ &+W{{d}_{i}}(H-{{m}_{hj}})(H-{{m}_{hj}}-1)]\end{split}$ (11)
@@
[…]
@@
Appendix B Training Details
Training and Testing Datasets. We curated a dataset containing 1.73 million high-resolution, high-quality images, each with an aesthetic score of 5.5 or higher. Following the production methodology outlined in ControlNet (Zhang et al. 2023), we generated various types of conditional images, including HED, Canny, Depth, and Segment. For the quantitative experiments, the model was trained at a resolution of 512 and evaluated on the COCO validation dataset (Lin et al. 2014), consisting of 5000 images. For qualitative visual evaluation, the model was trained at a resolution of 1024 and tested on a high-quality, high-resolution test set of 1000 images. To ensure a fair evaluation, all experiments were conducted with a training of 5 epochs on 16 NVIDIA A100 GPUs, using identical settings across all trials.
Additional experimental results
More exploration on Relevance
@@ Appendix B Training Details
More ablation results
To further investigate the effectiveness of the proposed modules, we conduct additional ablation studies focusing on two aspects: the performance of the TDSM and the impact of randomly selecting control positions. The detailed results are summarized in Table 5.
Table 5(a) illustrates that the random shuffle mechanism embedded in TDSM significantly enhances the interaction range within the model, facilitating more efficient extraction and propagation of control signals. Notably, replacing TDSM with conventional self-attention not only increases computational overhead but also degrades the generative performance. Furthermore, since the RGLC block simplifies the original PixArt Transformer block, we compare Setting 3, which retains a full feed-forward network and cross-attention, with the final RelaCtrl configuration. Despite a substantial increase in both parameter count and computational complexity, the fully structured Transformer block still underperforms compared to RelaCtrl, highlighting the superior efficiency and design of TDSM.
In addition, we include ablation results where 13 control blocks are randomly selected, as shown in Table 5(b). Compared to the ControlNet-top13 and Relevance-top13 settings, random selection results in noticeable degradation in both control accuracy and visual fidelity. This observation supports our previous findings: control relevance is more prominent in the early and middle layers of the DiT backbone, whereas random selection tends to distribute control modules more uniformly across the network (as illustrated in Fig. 1 of the manuscript). Because modules placed in deeper layers are less effective at integrating control signals, their contribution to overall performance is diminished.
Table 5: Supplementary ablation study results.
(a) TDSM Ablation Results
[…]
w/o Shuffle 97.29 21.63 15.70%
TDSM to SA 96.37 21.32 30.70%
RGLC to Copy 95.57 21.28 84.6%
RelaCtrl 94.04 20.34 15.3%
(b) Random Control Module Placement
[…]
Random-13 96.51 23.74 100%
ControlNet-top13 96.26 21.38 100%
Relevance-top13 94.57 20.31 100%
Experiment on Flux
@@
In this section, we apply RelaCtrl to the latest Flux.1-dev with 12 billion parameters for the experiment. The comparison methods include OminiControl (Tan et al. 2024) and Flux ControlNet, which is implemented based on the Diffusers library. Flux ControlNet utilizes a standard configuration consisting of a double-stream layer with the first four replication layers, alongside a single-stream layer with the first ten replication layers.
Our goal in integrating RelaCtrl into Flux is to transfer as much of the original architectural insight and design as possible, thereby evaluating the generalization ability of the proposed method. To this end, we proportionally scale the control positions from PixArt to match the architectural layout of Flux. Notably, in the double-stream layers, the control is applied only to the image branch; in the single-stream layers, the control signals are injected exclusively into the image tokens.
Table 6 presents the performance and efficiency comparison of various methods. GFLOPs are calculated at an image resolution of 512, while inference time is measured using 30 DDIM sampling steps at the same resolution. RelaCtrl demonstrates superior performance across multiple quality metrics related to image generation. Although OminiControl introduces only a modest increase in parameter count, it results in a substantial rise in computational cost and inference time, thereby severely constraining its practical applicability. Flux-ControlNet exhibits moderate overhead across various efficiency indicators.
In contrast, RelaCtrl for Flux achieves a notable reduction in computational complexity and inference latency, striking a favorable balance between efficiency and model size. As illustrated in Fig. 7, RelaCtrl yields visually precise and detailed control outcomes, while also offering significant computational advantages. These results collectively demonstrate the effectiveness, efficiency, and generalizability of RelaCtrl across different DiT-based architectures.
Table 6: Quantitative comparison results based on the Flux.1-dev. The best results are in bold, and the second-best results are underlined.
Setting FID $\downarrow$ HDD $\downarrow$
CLIP-
Score $\uparrow$
CLIP-
Ae $\uparrow$
Parameters (M)
Complexity
Inference
Flux.1-dev - - - - 11901.39 9925.78 4.78
+ControlNet-canny 17.61 87.26 0.2813 5.2871 +2952.65 +2578.33 +0.79
+OminiControl-canny 17.19 87.38 0.2837 5.1591 +14.49 +6637.41 +3.11
+RelaCtrl-canny 16.86 86.87 0.2861 5.2039 +549.19 +495.03 +0.34
Figure 7: Visual comparison of RelaCtrl and OminiControl on Flux.1-dev.
Inference with Community Models
@@
We perform inference using the PixArt weights, which were fine-tuned with Lora. Although RelaCtrl has not previously been trained in these weights, it can still utilize them effectively. Fig. 8 showcases the model’s generated paint, oil, gufeng, and pixel-style images under the specified conditions.
@@
Efficiency across diverse platforms
To further evaluate the computational efficiency of RelaCtrl across diverse hardware platforms, we conduct comprehensive runtime benchmarks on three distinct devices: NVIDIA A100, NVIDIA RTX 3090, and Huawei Ascend-910B. The RTX 3090 represents a widely adopted consumer-grade GPU, whereas the A100 and Ascend-910B are high-performance, industrial-grade accelerators featuring fundamentally different architectures. Detailed benchmark results are reported in Table 7. Across all tested platforms and workloads, our method consistently demonstrates high runtime efficiency and stable performance. These results further underscore the robustness and hardware-agnostic efficiency of RelaCtrl, highlighting its practicality for deployment in both consumer-level and industrial-scale applications.
Table 7: Effectiveness evaluation across different hardware platforms.
Methods Base model Parameters Complexity
Inference (20 steps)
Memory
Pixart-a 611.15 M
542.56 on A100
266.78 on 3090
479.01 on 910B2
3.81 Sec on A100
9.42 Sec on 3090
5.92 Sec on 910B2
2114 MiB on A100
2130 MiB on 3090
2297 MiB on 910B2
+ControlNet Pixart-a +294.34 M
+270.57 on A100
+128.61 on 3090
+246.16 on 910B2
+0.51 Sec on A100
+0.59 Sec on 3090
+3.99 Sec on 910B2
+1694 MiB on A100
+1743 MiB on 3090
+1722 MiB on 910B2
+RelaCtrl +45.15M
+46.71 on A100
+28.50 on 3090
+40.65 on 910B2
+0.24 Sec on A100
+0.44 Sec on 3090
+1.24 Sec on 910B2
+395 MiB on A100
+582 MiB on 3090
+735 MiB on 910B2